Find the speed of sound in water which has a Young's Modulus of 2.1×10^-9 N/m2 at a temperature of 0 and a density of 1000kgm/m3


1
Expert's answer
2021-03-10T17:16:24-0500

Answer

The speed of sound in water at oC can be calculated as

"v =\sqrt{\frac{ (2.1\times10^{-9} ) }{(1000)}}

 \\ = 1.45\times10^{-6}m/s"

where

Ev= 2.1"\times10^{-9}"  (N/m2)

and

ρ =1000 (kg/m3)